# Information Loss under Coarse-Grained Partitions: A Discrete Framework for AI Ethics

### v6

@@ Abstract
As artificial intelligence (AI) systems become increasingly embedded in ethically sensitive domains such as education, healthcare, and transportation, the need to balance accuracy and interpretability in decision-making has become a central concern. Coarse Ethics (CE) motivates the use of coarse-grained evaluations, such as letter grades, as ethically appropriate under cognitive, institutional, and contextual constraints. However, CE has lacked a simple mathematical formalization that makes the informational cost of such coarse judgments explicit. This paper introduces coarse-grained partitions (CGPs) as a discrete, information-theoretic framework for modeling coarse evaluation. We treat an underlying score scale as a finite ordered set, represent coarse evaluation as a partition into grains together with an index assignment, and define the induced coarse-grained distribution by pushforward. To quantify the informational impact of coarse-graining, we propose categorical unification (CU), which constructs a canonical within-grain unification of the original distribution, and we measure information loss by the discrete Kullback–Leibler divergence between the original distribution and its CU-unification (KL-CU). We show that information loss is zero if and only if the original distribution is already uniform within each grain, clarifying precisely when coarse evaluation preserves all relevant information. We illustrate the framework with educational grading and explainable AI (XAI) settings, where coarse categories are often required for communication and accountability. By providing a tractable discrete formalization of coarse evaluation and a principled measure of information loss, CGPs offer a foundation for designing interpretable AI systems that balance comprehensibility, fairness, and informational integrity.
@@ 1.1 Background
From an ethical standpoint, Coarse Ethics (CE, [izumoweng2022]) proposes that coarse-grained assessments are not only necessary but also ethically justifiable in many decision-making contexts. For example, academic grading systems classify students into letter grades or grade point average (GPA), abstracting away precise score differences. Similarly, phrases like “Olivia and Noah are Harvard graduates” generalize individual performance details. In AI-driven decision-making, CE argues that the legitimacy of transforming fine-grained evaluations into coarse ones depends on the cognitive capacity and contextual needs of the audience [izumo2025].
1.2 Challenges in Coarse Ethics
Despite its conceptual appeal, CE faces two primary challenges:
  1. 1.
    Lack of Mathematical Rigor: While CE provides an ethical justification for coarse evaluations, it lacks a mathematically formalized framework. For instance, Izumo [izumo2025] proposes that an AI assistant should issue coarse warnings such as “Your driving is dangerous” rather than continuously recalculating and displaying the optimal speed. However, the study does not specify the exact computational method to determine when and how such warnings should be issued. Without a formalized structure, these decisions may be inconsistent, raising concerns about reliability in AI-driven risk assessment.
  2. 2.
    Ambiguity in Categorization: Prior research has introduced a three-tiered explanatory approach—detailed explanations for experts, coarse explanations for general users, and even coarser explanations for individuals with cognitive limitations [izumo2025]. However, this classification is overly simplistic and lacks adaptability, failing to account for the diverse range of cognitive and contextual factors in human decision-making.
Addressing these limitations is essential, particularly as AI advances toward artificial general intelligence (AGI). Strengthening the mathematical foundation of CE and refining its categorization mechanisms can contribute to the development of more robust, context-aware AI explanations.
@@ 1.3 Methodology
Several prior studies share conceptual similarities with CE. For example, He et al. [he2024analogies] explored how analogies can assist laypeople in AI-assisted decision-making. Their work emphasizes the cognitive aspects of explanation reception and user comprehension. While our study shares the objective of improving human-AI interaction, we focus on the mathematical quantification of information loss in coarse evaluations rather than the psychological mechanisms of explainability. Peters and Carman [peters2024cultural] examined the influence of cultural bias in XAI research, demonstrating how users from different cultural backgrounds interpret AI explanations differently. Their study focuses on the subjective perception of AI explanations rather than developing a mathematical framework for evaluation. In contrast, our work formalizes coarse-grained evaluations using set theory, providing a structured approach to managing granularity in AI-driven decision-making.
This paper proposes coarse-grained partitions (CGPs) as a novel set-theoretic approach to resolving these issues. Set theory has been extensively applied to human decision-making, forming the basis of fuzzy set theory [zadeh1965fuzzy], rough set theory [pawlak1982rough], and soft set theory [molodtsov1999soft]. These approaches primarily focus on modeling real-world uncertainty through mathematical representations.
In contrast, CGPs shift the focus from enhancing AI’s ability to analyze the world to helping humans extract and interpret AI-generated insights. While previous works aimed at improving AI’s analytical capabilities with set theory [10.5120/8924-2996, 9015977, dalkilic2022novel], CGPs are based on the premise that AI models now outperform human experts in specific domains, as demonstrated by recent studies on AI surpassing human-level performance in standardized exams [bicknell2024chatgpt, katz2023gpt4]. Consequently, the key question is no longer how AI perceives reality, but rather how humans can efficiently understand and interact with AI-driven decisions.
This perspective is particularly relevant as Large Language Models (LLMs) and domain-specific AI systems continue to exceed human performance in complex tasks [vaswani2017attention, radford2018improving, brown2020language, chowdhery2022palm]. As AI models continue to outperform human experts in specialized tasks, discussions surrounding AGI expectations and ethical considerations have intensified [morris2024levelsagioperationalizingprogress]. By establishing a formal mathematical foundation for coarse evaluations, CGPs offer a critical framework for addressing these challenges and redefining the evolving relationship between AI and human cognition.
2 Coarse-Grained Partitions
Throughout the paper, we fix a finite totally ordered set $(U,\leq)$, which we call the underlying scale. In applications, one may take $U=\{0,1,\dots,100\}$ with the natural order. We equip $U$ with the discrete $\sigma$-algebra $2^{U}$.
@@ 2.1 Coarse-Grained Partitions
Let $(U,\leq)$ be a finite totally ordered set, and let $\mathscr{I}(U,\leq)$ denote the family of all nonempty order-convex (interval) subsets of $U$. A coarse-grained partition of $(U,\leq)$ is a partition object $\pi$ equipped with a finite index set $I_{\pi}$ and a family of grains
@@
Here $\pi$ denotes the partition object, while $\mathfrak{G}_{\pi}$ denotes its underlying family of grains.
Remark 1 (On generality).
The discrete ordered framework above is sufficient for the main results of this paper and matches typical applications such as grading scales. A measure-theoretic extension to general measurable spaces can be developed separately (e.g. by replacing sums with integrals), and may be placed in an appendix if needed.
Remark 2 (Ordering grain families).
Remark 3 (No internal order within grains).
In our constructions, each grain $G\in\mathfrak{G}_{\pi}$ is treated as an unordered set: only membership $u\in G$ and the cardinality $|G|$ are used. In particular, the CU construction assigns probabilities uniformly within each grain, so no internal ordering of elements inside a grain is required in this paper. By contrast, in extensions where one selects a representative value for each grain—for instance, taking $\min(G)$ as the representative—the ambient order on $U$ (and hence the induced order within $G$) becomes relevant.
@@
The collection of (interval) coarse-grained partitions of $U$ is denoted by $\mathcal{P}_{\textrm{int}}(U)$ in this paper. In particular, for the example $U_{1}=\{1,2,3\}$ above, we have
$\mathcal{P}_{\textrm{int}}(U_{1})=\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\},$
where each partition $\pi_{k}$ is represented by its grain family $\mathfrak{G}_{\pi_{k}}$ as listed in the example. Accordingly, the collection of grain families induced by interval partitions is
$\mathscr{G}_{\textrm{int}}(U_{1})=\{\mathfrak{G}_{\pi_{1}},\mathfrak{G}_{\pi_{2}},\mathfrak{G}_{\pi_{3}},\mathfrak{G}_{\pi_{4}}\}.$
@@
3 Coarse Assignment and Category Assignment
3.1 Coarse Assignment
CE studies how a fine-grained evaluation on a scale $U$ is mapped into a coarser categorization. Given a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, the coarse categories are conceptually given by the grains $G_{\pi,i}\in\mathfrak{G}_{\pi}$. For technical convenience (measurability and pushforward constructions), we represent these categories by their labels $I_{\pi}$ and fix a coarse assignment map
@@ 3.1 Coarse Assignment
Definition 2 (Index set and coarse assignment map).
Let $(U,\leq)$ be a finite totally ordered set, equipped with the discrete $\sigma$-algebra $2^{U}$. Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ (cf. Def. 1). Define the index set by
$I_{\pi},$
equipped with the discrete $\sigma$-algebra $2^{I_{\pi}}$. The coarse assignment map associated with $\pi$ is the map $q_{\pi}:U\to I_{\pi}$ defined by
@@
Example 2 (A coarse assignment map).
@@
Define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by $q_{\pi}(u)=i$ iff $u\in G_{\pi,i}$. Then
@@
Remark 4 (Measurability in the finite/discrete setting).
Since we work with the discrete $\sigma$-algebras $2^{U}$ and $2^{I_{\pi}}$ on finite sets, every map between these spaces is measurable. In particular, the coarse assignment map $q_{\pi}:U\to I_{\pi}$ is automatically measurable.
3.2 Object-to-Category Map
In practice, an evaluation is always performed of some collection of objects. Therefore, it is not enough to specify only the relationship between a fine-grained scale and its coarse-grained counterpart. One must also fix (i) how objects are mapped to fine-grained scores, and (ii) how those scores are subsequently mapped to coarse categories. In other words, we must specify both the object-to-score map and the induced object-to-category assignment.
@@ 3.2 Object-to-Category Map
Let $O$ be a nonempty set of objects and let $f:O\to U$ be an object-to-score map. Fix a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ and coarse assignment map $q_{\pi}:U\to I_{\pi}$. The object-to-category map induced by $f$ and $\pi$ is the composition
@@ 4.1 Probability Measures
Since normative assessments are typically made over a finite set of cases in practice, this section restricts attention to the finite setting. We model both objects and scores as finite sets and represent empirical frequencies by discrete probability distributions. This choice keeps the information-loss analysis entirely within the standard discrete (Shannon/KL) formalism, without comparing probability measures of different types (e.g., continuous vs. atomic).
@@
Let $U$ be a finite set and equip it with the power set $\Sigma:=2^{U}$. Let $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ be a finite coarse-grained partition of $U$, and let $P_{U}$ be a probability distribution on $(U,2^{U})$. With $q_{\pi}$ as in Definition 2, define the coarse-grained probability distribution $P_{\pi}$ on $(I_{\pi},2^{I_{\pi}})$ by the pushforward
$P_{\pi}:=P_{U}\circ q_{\pi}^{-1}.$
Equivalently, for each $i\in I_{\pi}$,
$P_{\pi}(\{i\})=P_{U}(G_{\pi,i}).$
Remark 6 (The coarse $\sigma$-algebra generated by $\pi$).
The partition $\pi$ induces the $\sigma$-algebra
$\Sigma_{\pi}:=\sigma(\pi)=\Bigl\{\bigcup_{i\in J}G_{\pi,i}\;:\;J\subseteq I_{\pi}\Bigr\}\subseteq\Sigma.$
Moreover, $q_{\pi}$ is $(\Sigma,\;2^{I_{\pi}})$-measurable and $\Sigma_{\pi}=q_{\pi}^{-1}(2^{I_{\pi}})$.
The space $O$ represents the underlying domain of objects (e.g., test scores), equipped with a probability distribution $\mathbb{P}$ that captures how likely different objects are to occur. The map $f:O\to U$ assigns each object in $O$ to a category in $U$. Definition 4 defines the induced probability distribution on $U$ by pushing forward $\mathbb{P}$ through $f$: for any measurable set of categories $A\subseteq U$, the probability that the assigned category lies in $A$ is defined as the probability mass of all objects in $O$ that are mapped into $A$, i.e.,
$\mathbb{P}_{U}(A)=\mathbb{P}(f^{-1}(A)).$
In other words, $\mathbb{P}_{U}$ is the distribution of the categorized outcome $f(o)$ when $o$ is drawn according to $\mathbb{P}$.
Example 3 (Pushforward to a win/lose outcome).
As a running toy example, consider a dice game in which a larger face value yields a higher score. We take the underlying scale to be
$U:=\{1,2,3,4,5,6\}\quad\text{equipped with the natural total order }1\leq 2\leq\cdots\leq 6.$
Let $O:=\{a,b,c,d,e\}$ be a set of five players (or five trials), and define an outcome map $f:O\to U$ by
$f(a)=1,\quad f(b)=1,\quad f(c)=3,\quad f(d)=5,\quad f(e)=6.$
The induced empirical distribution $\hat{P}\in\Delta(U)$ is defined by
$\hat{P}(u):=\frac{|\{o\in O:f(o)=u\}|}{|O|}\qquad(u\in U).$
Hence,
$\hat{P}(1)=\frac{2}{5},\ \hat{P}(2)=0,\ \hat{P}(3)=\frac{1}{5},\ \hat{P}(4)=0,\ \hat{P}(5)=\frac{1}{5},\ \hat{P}(6)=\frac{1}{5}.$
Next, define a two-grain partition $\pi$ of $U=\{1,2,3,4,5,6\}$ by the grain family
$G_{\pi,1}:=\{1,2,3\},\qquad G_{\pi,2}:=\{4,5,6\},$
and denote the associated grain family by
$\mathfrak{G}_{\pi}:=\{G_{\pi,1},G_{\pi,2}\}.$
(For readability, we may refer to $G_{\pi,1}$ as the Lose-grain and $G_{\pi,2}$ as the Win-grain.)
Let $I_{\pi}:=\{1,2\}$ and define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by
$q_{\pi}(u)=\begin{cases}1,&u\in G_{\pi,1}=\{1,2,3\},\\ 2,&u\in G_{\pi,2}=\{4,5,6\}.\end{cases}$
Introduce the grain-label map $\ell:I_{\pi}\to\mathfrak{G}_{\pi}$ by
$\ell(1)=G_{\pi,1},\qquad\ell(2)=G_{\pi,2}.$
Then $\ell\circ q_{\pi}:U\to\mathfrak{G}_{\pi}$ maps each outcome $u\in U$ to its grain, and hence
$\ell\circ q_{\pi}\circ f:O\to\mathfrak{G}_{\pi}.$
The pushforward distribution on grains, $\hat{P}_{\mathfrak{G}}:=(\ell\circ q_{\pi})_{*}\hat{P}\in\Delta(\mathfrak{G}_{\pi})$, is
$\hat{P}_{\mathfrak{G}}(G):=\sum_{u\in(\ell\circ q_{\pi})^{-1}(G)}\hat{P}(u)\qquad(G\in\mathfrak{G}_{\pi}),$
so that
$\hat{P}_{\mathfrak{G}}(G_{\pi,1})=\hat{P}(1)+\hat{P}(2)+\hat{P}(3)=\frac{3}{5},\qquad\hat{P}_{\mathfrak{G}}(G_{\pi,2})=\hat{P}(4)+\hat{P}(5)+\hat{P}(6)=\frac{2}{5}.$
4.2 Kullback–Leibler Divergence
$D_{\text{KL}}(P\parallel Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (1)
To illustrate the motivation behind coarse ethics, consider now that we are a small classroom teacher and give a math test to ten students. The results appear in Table 1, Figure 1.
@@ 4.2 Kullback–Leibler Divergence
Define the score map $f:O\to U$ from Table 1 by
@@
From the score list, the counts are
$\#\{u=0\}=1,\ \#\{u=2\}=1,\ \#\{u=3\}=3,\ \#\{u=4\}=1,$
$\#\{u=5\}=2,\ \#\{u=6\}=1,\ \#\{u=10\}=1,$
and all other scores occur 0 times. Hence
@@
Define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by
@@
Define the pushforward distribution on labels by
$\hat{P}_{L}:=(\ell\circ q_{\pi})_{*}\hat{P}\in\Delta(L),\qquad\hat{P}_{L}(\lambda):=\sum_{u\in(\ell\circ q_{\pi})^{-1}(\lambda)}\hat{P}(u)\quad(\lambda\in L).$
@@
We compare the probability distributions between the original grading system (ranging from 0 to 10) and the binary pass/fail assessment. More intuitively, if we inform ten students only of the pass/fail criteria and their respective pass or fail status, without disclosing their specific scores—or if we later lose the list of scores and can only recall the pass/fail criteria and each student’s pass or fail status—our task is to determine how much information has been lost.

### v7

# Quantifying Information Loss under Coarse-Grained Partitions: A Discrete Framework for Explainable Artificial Intelligence

@@ Abstract
As artificial intelligence (AI) systems are increasingly used in ethically sensitive domains such as education, healthcare, and transportation, balancing accuracy and interpretability has become a central concern. Coarse ethics (CE) motivates coarse-grained evaluations under cognitive, institutional, and contextual constraints, but it still lacks a simple mathematical formalization of admissible coarse-graining and its informational consequences. This paper introduces coarse-grained partitions (CGPs) as a discrete framework for modeling coarse evaluation on a finite totally ordered score scale. A CGP represents coarse evaluation as a partition into grains with an index assignment, and induces a coarse-grained distribution by pushforward. To compare admissible coarse-grainings, we introduce categorical unification (CU), which constructs a canonical fine-scale reconstruction from the coarse representation under minimal assumptions. On this basis, we define a KL-based measure of information loss, $D_{\mathrm{KL\text{-}CU}}$, as the divergence between the original fine-grained distribution and its CU-based reconstruction. We prove that $D_{\mathrm{KL\text{-}CU}}=0$ if and only if the original distribution is already uniform within each grain. This shows that zero loss, in the sense of the proposed measure, is a highly exceptional limiting case rather than a realistic benchmark for ordinary evaluative practice. We also show that the framework leads naturally to an optimization problem for comparing alternative admissible CGPs. Applications to educational grading and explainable AI (XAI) illustrate how the framework clarifies trade-offs among informational fidelity, interpretability, and coarsening cost.
@@ 1.1 Background
Recent research on AI-assisted decision-making has increasingly emphasized that explanation and evaluation are not purely technical matters, but also ethical ones. In particular, a growing body of work has examined how AI outputs should be presented to human users in ways that are cognitively manageable, socially appropriate, and normatively defensible. Several prior studies are relevant in this respect. For example, He et al. [he2024analogies] explored how analogies can assist laypeople in AI-assisted decision-making, highlighting the cognitive conditions under which explanations become understandable to non-expert users. Peters and Carman [peters2024cultural] examined the influence of cultural bias in XAI research, showing that users from different cultural backgrounds may interpret AI explanations differently. These studies underscore that the form and granularity of AI explanations are not neutral, but depend on the capacities, expectations, and interpretive contexts of their audiences.
From this broader ethical perspective, coarse ethics (CE, [izumoweng2022], [izumo2025]) offers a particularly suggestive framework. CE proposes that coarse-grained assessments are not only practically unavoidable but also ethically justifiable in many decision-making contexts. For example, academic grading systems classify students into letter grades or grade point averages (GPAs), thereby abstracting away precise score differences, and expressions such as “Olivia and Noah are Harvard graduates” similarly suppress finer distinctions in individual performance. In AI-driven decision-making, CE argues that the legitimacy of transforming fine-grained evaluations into coarse ones depends on the cognitive capacity and contextual needs of the audience [izumo2025].
The present paper focuses on this CE perspective. Despite its conceptual appeal, CE faces a central challenge: it lacks a mathematically rigorous structure, and this weakens the usefulness of the very notion of coarse-grained normative evaluation. Izumo and Weng [izumoweng2022] identify two requirements for coarsening a fine-grained evaluation: first, that the coarse-grained evaluation should sufficiently cover the objects evaluated in the original one (for example, if Olivia is evaluated in the original assessment, she should not disappear from the coarse-grained version); and second, that the coarse-grained evaluation should not reverse the value ordering embodied in the fine-grained one (for example, if Olivia is rated higher than Noah in the original evaluation, the coarsened evaluation should not end up ranking Noah above Olivia). However, as the authors themselves acknowledge, these conditions do not suffice to determine a unique coarse-grained evaluation. As a result, multiple distinct coarse evaluations satisfying those conditions may arise from a single fine-grained one. This problem stems from the fact that the previous literature does not specify the mathematical process by which coarsening is carried out.
@@ 1.2 Methodology
This paper proposes coarse-grained partitions (CGPs) as a novel set-theoretic framework for addressing these issues. Set-theoretic approaches have long been used in the study of human decision-making, most notably in fuzzy set theory [zadeh1965fuzzy], rough set theory [pawlak1982rough], and soft set theory [molodtsov1999soft]. These approaches were primarily developed to represent uncertainty, vagueness, and incomplete information in the world as mathematically tractable structures. CGPs differ from these approaches in their point of departure. Rather than improving AI’s ability to model external reality, CGPs are designed to analyze how fine-grained evaluations can be transformed into coarser forms that remain usable for human agents. In this sense, the problem is no longer how AI perceives the world, but how human beings can understand, interpret, and act upon AI-generated evaluations. This perspective becomes especially important in domains where AI systems now outperform human experts in narrowly defined tasks, including standardized testing and other forms of specialized assessment [bicknell2024chatgpt, katz2023gpt4]. As AI systems become increasingly capable, the challenge of translating their internal evaluations into forms intelligible to human users becomes correspondingly more urgent.
From this perspective, CGPs provide a mathematical framework for studying coarse-grained evaluation itself. We formalize the transformation from fine-grained to coarse-grained evaluation by introducing partitions of a finite discrete totally ordered score scale. We then compare the probability distributions before and after coarsening, define a canonical reconstruction of the fine-grained distribution through categorical unification, and measure information loss by KL divergence [kullbackleibler1951] with categorical unification (CU), i.e., $D_{\mathrm{{KL\text{-}CU}}}$. This allows us to analyze, in precise terms, when coarse-graining preserves information, when it necessarily destroys it, and how alternative coarse-graining schemes may be compared. More precisely, the role of CGPs in this paper is twofold. First, as interval partitions of a finite totally ordered scale, they provide a mathematically explicit class of coarse-grainings that satisfies the structural requirements emphasized in CE, namely coverage and order preservation. Second, since these requirements alone do not determine a unique coarse-graining, we introduce $D_{\mathrm{{KL\text{-}CU}}}$ not as a replacement for them, but as an additional criterion for comparing admissible CGPs in terms of informational loss.
The structure of the paper is as follows. Section 2 introduces the basic framework of coarse-grained partitions on finite discrete totally ordered sets. Section 3 formalizes the object-to-category map and the induced probability distributions associated with coarse-grained evaluation. Section 4 defines CU and the KL-based measure of information loss, and proves the zero-information-loss theorem. Section 5 examines the optimization problem of coarse-graining design and discusses its significance for explainable AI and related decision-making contexts. Section 6 concludes.
2 Coarse-Grained Evaluation
@@ 2.1 Coarse-Grained Partitions
In many institutional settings, normative evaluations are implemented on finite, discrete score sets. For instance, school test scores are recorded as bounded integers rather than real numbers. In this paper, we take such discretized scores as our main motivating case. Accordingly, it suffices to model the fine-grained score domain as an arbitrary finite totally ordered set (e.g., a finite subset of $\mathbb{N}$). Throughout the paper, we fix a finite totally ordered set $(U,\leq)$, which we call the underlying scale. In applications, one may take $U=\{0,1,\dots,100\}$ with the natural order. We equip $U$ with the discrete $\sigma$-algebra $2^{U}$.
What must be discussed rigorously, when we speak of coarsening a fine-grained evaluation, is how the resulting coarse-graining should be represented as a mathematical object. Here we formalize it as a coarse-grained partition of $U$, which will later induce a score-to-category map.
@@
Given the underlying scale $(U,\leq)$, let $\mathscr{I}(U,\leq)$ denote the family of all nonempty intervals of $U$. A coarse-grained partition of $(U,\leq)$ is a partition object $\pi$ equipped with a finite index set $I_{\pi}$ and a family of grains
@@
Here $\pi$ denotes the structure specifying how $U$ is partitioned, while $\mathfrak{G}_{\pi}$ denotes its family of grains.
@@
Remark 1 (Ordering grain families).
@@
3 Object-to-Category Map
3.1 Definition
CE studies how a fine-grained evaluation on a scale $U$ is mapped into a coarser categorization. In practice, an evaluation is always performed of some collection of objects. For example, the evaluation “Emily is in rank A because she scored 90 on a math test” can be viewed as a two-step mapping: Emily (as an object) is first mapped to a fine-grained numerical score (90 points), and this score is then mapped to a coarse category (rank A). Accordingly, we need a map from each object to a category (an object-to-category map); however, this map is the composition of two maps: one from the object to a fine-grained score (an object-to-score map), and another from that fine-grained score to a coarse category (a score-to-category map).
We begin by defining the score-to-category map. Given a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, the coarse categories are conceptually given by the grains $G_{\pi,i}\in\mathfrak{G}_{\pi}$. For technical convenience (measurability and pushforward constructions), we represent these categories by their labels $I_{\pi}$ and fix a score-to-category map
@@ 3.1 Definition
Definition 2 (Score-to-category map).
Let $(U,\leq)$ be a finite totally ordered set, equipped with the discrete $\sigma$-algebra $2^{U}$. Let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ (cf. Def. 1). Define the index set $I_{\pi}$ as the label set appearing in $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, equipped with the discrete $\sigma$-algebra $2^{I_{\pi}}$. The score-to-category map associated with $\pi$ is the map $q_{\pi}:U\to I_{\pi}$ defined by
@@
Example 2 (A score-to-category map).
@@
Define the score-to-category map $q_{\pi}:U\to I_{\pi}$ by $q_{\pi}(u)=i$ iff $u\in G_{\pi,i}$. Then
@@
Let $O$ be a nonempty set of objects and let $f:O\to U$ be an object-to-score map. Fix a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ and score-to-category map $q_{\pi}:U\to I_{\pi}$. The object-to-category map induced by $f$ and $\pi$ is the composition
@@ 4.1 Probability Measures
Since the previous section defined the evaluation objects and the maps $f:O\to U$ (from objects to fine-grained scores) and $q_{\pi}:U\to I_{\pi}$ (from fine-grained scores to category labels), we now compare the induced distributions on the fine-grained score scale $U$ and on the label set $I_{\pi}$. Concretely, given a probability distribution on $O$, we obtain the score distribution on $U$ as the pushforward along $f$, and the coarse distribution on $I_{\pi}$ as the pushforward along $q_{\pi}\circ f$.
@@
Given the underlying scale $(U,\leq)$, let $\pi$ be a coarse-grained partition of $(U,\leq)$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, and let $P_{U}$ be a probability distribution on $(U,2^{U})$. With $q_{\pi}$ as in Definition 2, we define the coarse-grained probability distribution $P_{\pi}$ on $(I_{\pi},2^{I_{\pi}})$ as follows. First, for each label $i\in I_{\pi}$, set
$P_{\pi}(\{i\}):=P_{U}(G_{\pi,i}).$
We then extend this to arbitrary subsets $A\subseteq I_{\pi}$ by additivity, defining
$P_{\pi}(A):=\sum_{i\in A}P_{\pi}(\{i\})=\sum_{i\in A}P_{U}(G_{\pi,i})=P_{U}\!\left(\bigcup_{i\in A}G_{\pi,i}\right).$
Equivalently, for every $A\subseteq I_{\pi}$,
$P_{\pi}(A)=P_{U}\bigl(q_{\pi}^{-1}(A)\bigr),$
that is, $P_{\pi}$ is the pushforward of $P_{U}$ along $q_{\pi}$.
Example 3 (Pushforward to a pass/fail outcome).
As a running example, consider now that we are a small classroom teacher and give a math test to ten students. The results appear in Table 1, Figure 1.
@@
Define the object-to-score map $f:O\to U$ from Table 1 by
@@
Hence
@@
Define the score-to-category map $q_{\pi}:U\to I_{\pi}$ by
@@
Define the pushforward distribution on labels $\hat{P}_{L}$ on $(L,2^{L})$ by
$\hat{P}_{L}(A):=\hat{P}\bigl((\ell\circ q_{\pi})^{-1}(A)\bigr)\qquad(A\subseteq L).$
In particular, for a singleton $A=\{\lambda\}$ with $\lambda\in L$, we have
$\hat{P}_{L}(\{\lambda\})=\hat{P}\bigl((\ell\circ q_{\pi})^{-1}(\{\lambda\})\bigr)=\sum_{u\in(\ell\circ q_{\pi})^{-1}(\{\lambda\})}\hat{P}(u).$
@@
4.2 Kullback–Leibler Divergence
4.2.1 Setup
$D_{\text{KL}}(P\parallel Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (1)
Using Example 3, we compare the distributions induced by the original 0–10 scoring scheme and by its binary pass/fail coarsening. Intuitively, if only pass/fail outcomes are retained (and the exact scores are discarded), we would like to quantify how much information is lost.